OpenStudy (buttermequeasy):

A frequency distribution is shown below. The number of dogs per household in a small town. Dogs,Households 0,1239 1,422 2,169 3,41 4,27 5,17 a. use the frequency distribution to construct a probability distribution B. find the mean of the probability distribution c. find the variance d. find the standard deviation e. interpret the results in the context of a real life situation household on average has 0.9 dog with a standard deviation of 0.9 dog. B. A household on average has 0.6 dog with a standard deviation of 0.9 dog.

OpenStudy (buttermequeasy):

A. household on average has 0.9 dog with a standard deviation of 0.9 dog. B. A household on average has 0.6 dog with a standard deviation of 0.9 dog. C. A household on average has 0.6 dog with a standard deviation of 17 dog. D. A household on average has 0.9 dog with a standard deviation of 0.6 dog

OpenStudy (buttermequeasy):

please help omg my grade is riding on this :-(

OpenStudy (campbell_st):

so the probability distribution is just the relative frequency 1st add the total number of households in the survey so add a column to the table and then for 0 dogs its 1239/Total households write the answer as a decimal, maybe 3 decimal places same for 1 dog 422/ total households.... repeat the process for the remainder of the table. then you should be able to do the rest...

OpenStudy (buttermequeasy):

i already did that part i neeed the variance and the standard deviation and the probability distribution :-(

OpenStudy (campbell_st):

well isn't the standard deviation \[SD = \sqrt{n} \times p \times (1 - p)\] perhaps this may help http://www.mathsisfun.com/data/random-variables-mean-variance.html

OpenStudy (buttermequeasy):

forget it. thanks i guess.